# Modal Verification Patterns for Systems

### v2

@@ 2.2.3 Reference-Counting Memory Reclamation
To reclaim the allocated memory location that is no longer being accessed, many systems use a well-known method called reference-counting in which each memory allocation includes not only its data but also a counter of how many pointers to that memory exist in the rest of the program. Any time a new reference to the allocated memory is created, the counter is incremented, and every time an existing reference is destroyed (by overwriting or dropping) the counter is decremented. [^1] This may be done using library functions for a reference counting system, or the operations may be inserted by a compiler. These reference counts are a sound (but incomplete) way to approximate reachability: if a decrement of a reference count decrements the count to 0, no other references to that memory exist in the program, and the memory can be reclaimed.
@@
2.2.4 Weak Memory Models
Multi-core and multi-processor systems do not directly write every memory update back to memory (or even a coherent hardware cache) individually, as this would be extremely slow. Instead, they effectively buffer reads and writes to and from memory locally within each CPU core [38] [^2] This is a layer below the better-known L1 cache, which is coherent across all CPUs in the system. so that operations can be partially reordered and batched to improve performance. For most code — where the different CPUs are doing independent work — this is transparent to software. But fine-grained concurrent systems code which has multiple CPU cores simultaneously coordinating access to shared memory must reason about these (effective) bufferings and reorderings, whose details vary across CPU architectures.
Most of this coordination comes through reasoning about what will be true after local operations are published to memory, or what will be true after remote operations become visible locally [45, 10, 13]. Each of these flavors of contingent reasoning depend on use of certain fence or memory barrier operations, which block execution of a CPU until either all incoming or outgoing memory transactions are complete.
3 Contingency Decomposition of a System
@@
Shallow Modalities in the Iris Separation Logic
[…]
@@
This definition also reveals a subtlety about the denotation of assertions $\phi$ — they are effectively predicates of some kind, functions from some possible world into some logical algebra such as a BI-algebra [37]. Taking this interpretation literally gives rise to a convenient way to embed new modalities in Iris in a way well-suited to many use cases. If $B$ is a type with a BI-algebra structure, then any space of predicate functions (e.g., $\mathsf{val}\rightarrow B$ as a simplification) also carries a BI-algebra structure, lifted from $B$, often called the pointwise lifting of the algebra. A number of projects for for systems software verification now take advantage of this, explicitly or implicitly lifting the BI-algebra structure from Iris itself to Iris predicates used as assertions within modalities. Connecting this to Iris’s typeclasses for interactive proofs [30, 32] then permits working within the embedded logic as in Iris.
@@
In addition, the modal definitions themselves may represent ownership of resources in a substructural setting [41, 15, 29]. This allows modalities to simultaneously represent ownership of certain resources used for interpreting the modality itself, as well as whatever resource ownership is implied by the modal argument.
3.1 Decomposing a System into its Constituents Contingently
Table 1: Modal Decomposition of Program-Logics.
@@ 3.1 Decomposing a System into its Constituents Contingently
Post-Crash [5, 6, 7] $\Diamond\;P$ $\ell\mapsto_{n}^{\overline{\gamma}}v$ Strong Crash Recovery
[…]
StackRegion  [47] $\overset{ICut^{n}}{\hookrightarrow}\;P$ $\framebox{n}\;\ell\mapsto v$ Strong Alloc and Return to/from stack
Memory-Fence [13, 14, 9] $\triangle_{\pi}$ and $\triangledown_{\pi}$ $\ell\mapsto v$ Weak Fence Acquire and Release
Address-Space [33] [r]P $\ell\mapsto v$ Weak Address-Space Switch
Ref-Count [48] @ P $\ell_{1}\mapsto v$ Weak Allocating, Dropping and Sharing a Reference
*The StackRegion Modality is an instance of NextGen (called the Independence Modality in [47]).
@@
3.2 Resource
Consider first the address-space abstraction in an OS kernel. An address-space of a process is a container of virtual addresses referencing data in memory. One would expect to have points-to assertion from separation logic to specify ownership of a memory reference pointing to some data. But that ownership is relative to a specific address space — a specific container. We tend to think directly about what is true in an address space, with the simplest piece being an association between a virtual address and the data it points to. We call the simplest piece, in this and other examples, the resource element:
Definition 1 (Resource Elements)
@@ 3.2 Resource
Definition 2 (Resource Context)
@@
Except for Post-Crash-Modality, one can think of the resource contexts in the first column in Table 1 as containers for the corresponding resource elements in the second column.
@@
When reasoning about stack frame contents, the resource element would be a stack-memory points-to assertion ( $\framebox{n}\;\ell\mapsto v$) indicating that a certain offset into stack region $n$ holds value $v$.
@@
For virtual memory management, a virtual-points-to ownership assertion pairing a virtual address ($\ell$) with data ($v$) in an address space is natural. A process’s address space with the root address r is an abstraction that is treated as a container for virtual address mappings, $\ell\mapsto v$;
@@
When considering weak memory models, we consider address-value mappings of thread-local views in C11 memory model [34]). Modalities $\Delta_{\pi}(P)$ and $\nabla_{\pi}(P)$ with $P$ represent the memory address-value mappings that held before or will hold after certain memory fence operations by thread $\pi$.
@@
When dealing with reference-counting APIs, we may care to specify reachability of memory nodes ($\ell\mapsto v$) in a certain context defined by a shared root address. A shared memory address $\ell$ can be the root of the graph that can be a container of memory nodes ($\ell_{1}\mapsto v$) that are reachable from the root $\ell$.
@@
The resource element of Post-Crash Modality is not obvious in Table 1, and needs a bit of explanation. Perennial, based on the Iris logic, has both disk-points-to assertions $d[\textsf{p}]\mapsto_{n}v$ (for a specific disk $d$) and in-memory points-to assertions $\ell\mapsto_{n}^{\overline{\gamma}}$. Perennial crash-recovery logic book-keeps resource names (can be thought of as logical variables) $\overline{\gamma}$ to identify which assertions (resource elements) remain valid after a crash — these assertions are only usable while the names in $\overline{\gamma}$ are valid, and a crash resets them, discarding assumptions about volatile state. A subtlety of the notion of a resource context is that, unlike the earlier examples, the context does not need to be a literal data structure. It can instead be (various forms of) a set of executions, as in the Post-Crash and NextGen modalities. The Post-Crash modality $\diamond$ expresses that the assertion $P$ will be true after a crash discards all unstable storage (i.e., RAM). This was the inspiration for the NextGen modality, which is in fact a framework for defining “after-$t$” modalities, where $t$ is an transformation of the global state. [^3] The transformations are subject to some technical constraints that are unrelated to our point here.
3.3 Nominals for Systems Resource Contexts
@@
Strong nominalization is the case where resource elements must explicitly include the identity of their intended context, while weak nominalization occurs when the resource elements implicitly pick up the relevant context from how they are used. The first three modalities in Table 1 are strongly nominalized, with the resource elements generally carrying identifiers of a specific modality usage. [^4] The Post-Crash modality does not look like this in the presentation here; technically the definition of $\diamond$ quantifies over names $\overline{\gamma}$ internally, dealing with sets of possible contexts. The last three are weakly nominalized.
@@
By contrast, weakly nominalized elements are more concise, but then make talking about multiple contexts together marginally more complex: changing which modality an assertion is used with drastically changes its meaning. In the case of the Ref-Count modality, @$_{\ell}(\ell_{1}\mapsto v)$ says that $\ell$ points to a reference count wrapping $\ell_{1}$, while placing the $\ell_{1}\mapsto v$ under a jump modality for a different location entails talking about a different region of memory.
@@
4 Conclusion
We have argued that there is a systematic pattern emerging in the use of modalities in the verification of low-level systems software, and explained how a range of modern work fits into this pattern. While we have focused our attention on public verification work carried out in Iris, it also captures (a slightly more organized version of) our own thinking in coming up with designs for specifying distributed systems [19], and in ongoing work using different modalities for different parts of a copy-on-write filesystem (e.g., for assertions true in different snapshots). So not only are the modalities observably present, but they are demonstrably useful for conducting proofs of systems software. We believe that these essential patterns in the designs of modal abstractions for systems verification constitute a fundamental concept when working out how to specify different kinds of systems code.

### v3

@@ 2.2.3. Reference Counting Memory Reclamation
To reclaim the allocated memory location that is no longer being accessed, many systems use a well-known method called reference counting in which each memory allocation includes not only its data but also a counter of how many pointers to that memory exist in the rest of the program. Any time a new reference to the allocated memory is created, the counter is incremented, and every time an existing reference is destroyed (by overwriting or dropping) the counter is decremented. [^1] This may be done using library functions for a reference counting system, or the operations may be inserted by a compiler. These reference counts are a sound (but incomplete) way to approximate reachability: if a decrement of a reference count decrements the count to 0, no other references to that memory exist in the program, and the memory can be reclaimed.
@@
2.2.4. Weak Memory Models
Multicore and multiprocessor systems do not directly write every memory update back to memory (or even a coherent hardware cache) individually, as this would be extremely slow. Instead, they effectively buffer reads and writes to and from memory locally within each CPU core (Owens et al., 2009) [^2] This is a layer below the better-known L1 cache, which is coherent across all CPUs in the system. so that operations can be partially reordered and batched to improve performance. For most code — where the different CPUs are doing independent work — this is transparent to software. But fine-grained concurrent systems code which has multiple CPU cores simultaneously coordinating access to shared memory must reason about these (effective) bufferings and reorderings, whose details vary across CPU architectures.
Most of this coordination comes through reasoning about what will be true after local operations are published to memory, or what will be true after remote operations become visible locally (Sieczkowski et al., 2015; Dang et al., 2022; Doko and Vafeiadis, 2016). Each of these flavors of contingent reasoning depend on use of certain fence or memory barrier operations, which block execution of a CPU until either all incoming or outgoing memory transactions are complete.
2.2.5. Actor Model
An alternative approach to addressing the shared memory concurrency complication, i.e., data races, is to use actors  (Hewitt et al., 1973; agha1986actors) by eliminating shared mutable state and enforcing updates on exclusive local mutable state via exchanging immutable messages among processes. This is a way of sequentializing the concurrent interleaved execution. Often correctness of one actor depends on knowledge about other actors. In a typical consensus algorithm, for example, an operation is committed if a quorum of nodes reach an agreement on it.
Therefore, as expected, the specification of an actor in the current view would refer to the other actors’ states and has to validate that the shared knowledge amongst actors is preserved  (Gordon, 2019).
3. Contingency Decomposition of a System
@@
3.1. Encoding Modalities in Program Logics
Program logics which would like to enjoy modal propositions need to have an underlying model that allows encoding of the propositions. A Kripke model can be roughly seen as a triple of set of worlds (e.g., program states), a binary relation between worlds (e.g., transition relations between states) and a state interpretation function. For a modality $M$, $M(\phi)$ is true in one state/world $w$ if states related to $w$ by the binary relation satisfy $P$. Whether this is all related states, or just one, or something else varies depending on what $M$ is meant to model, but the modality is always tied to the relation of interest. For multimodal logics there are multiple relations. For example, for Dynamic Logic, there is a modality for each program, and the relation for running program “command” (in old program logic parlance) $C$ is the relationship between states before and after executing $C$.
Encoding Hoare Triples
Following Pratt’s original representation (Pratt, 1976), we can encode a Hoare triple $\{P\}C\{Q\}$ (asserting that if $P$ holds in some state, then executing $C$ will yield a state where $Q$ holds) as $P\rightarrow[C](Q)$. However, for this to hold for a specific programming language and assertion language, the modal representation requires a semantics in terms of an underlying model.
Conventionally, a Kripke model with its constituent must be picked. Gordon provides a Kripke model for an actor-based hybrid modal logic (Gordon, 2019), in which the transition relation is indexed by a particular command and a particular actor, relating pre- and post-states of a specific actor referenced by a nominal executing the command. Gordon encodes this hybrid logic as a library in the Dafny program logic (Leino, 2010) with the help of two state invariants and rely-guarantee assertions checking the stability of invariants. Wagner et al. take a similar approach to the design of their separation logic and utilize a Kripke model for reference-counted memory reasoning principles (Wagner et al., 2024).
On the other hand, program logics such as Iris (Jung et al., 2018) utilizes Kripke models for their separation algebras (a collection of CMRAs as an algebraic cousin of a BI-algebra (O’Hearn and Pym, 1999)), modeling modalities for both heap extension and step indexing (Nakano, 2000). However in general, many modalities can be given proof-theoretic semantics directly within a logic. This is the case for Iris’s weakest precondition modality (which is defined within the Iris logic itself), as well as for a wide range of other modalities in the literature (Restall, 2002). These modalities are typically multi-modal, with a form of roughly $M~\overline{ctxt}~\phi\triangleq\ldots\rightarrow\phi~\overline{ctxt}$, such that the evaluation of the modality incorporates the $\overline{ctxt}$ arguments into the evaluation of $\phi$.
Shallow Modalities in Separation Logic
@@ 3.1. Encoding Modalities in Program Logics
In a substructural setting (Restall, 1993; Krebbers et al., 2017a) the implication $\rightarrow$ is typically replaced with substructural implication $\mathrel{-\mkern-6.0mu*}$. This definition also reveals a subtlety about the denotation of assertions $\phi$ — they are effectively predicates of some kind, functions from some possible world into some logical algebra such as a BI-algebra (O’Hearn and Pym, 1999). Taking this interpretation literally gives rise to a convenient way to embed new modalities in Iris in a way well-suited to many use cases. If $B$ is a type with a BI-algebra structure, then any space of predicate functions (e.g., $\mathsf{val}\rightarrow B$ as a simplification) also carries a BI-algebra structure, lifted from $B$, often called the pointwise lifting of the algebra. A number of projects for systems software verification (Chajed, 2022; Chajed et al., 2023, 2019; Vindum et al., 2025; Doko and Vafeiadis, 2016, 2017; Dang et al., 2019; Kuru and Gordon, 2025; Wagner et al., 2024) now take advantage of this, explicitly or implicitly lifting the BI-algebra structure from Iris itself to Iris predicates used as assertions within modalities. Connecting this to Iris’s typeclasses for interactive proofs (Krebbers et al., 2018, 2017b) then permits working within the embedded logic just as if working in Iris.
@@
In addition, the modal definitions themselves may represent ownership of resources in a substructural setting (Restall, 1993; Došen, 1992; Kamide, 2002). This allows modalities to simultaneously represent ownership of certain resources used for interpreting the modality itself, as well as whatever resource ownership is implied by the modal argument.
3.2. Decomposing a System into its Constituents Contingently
Table 1. Modal Decomposition of Program-Logics.
@@ 3.2. Decomposing a System into its Constituents Contingently
Post-Crash (Chajed, 2022; Chajed et al., 2023, 2019) $\Diamond\;P$ $\ell\mapsto_{n}^{\overline{\gamma}}v$ Strong Crash Recovery
NextGen (Vindum et al., 2025) $\overset{t}{\hookrightarrow}\;P$ Own (t(a)) Strong Determined Based on the Model
StackRegion  (Vindum et al., 2025) $\overset{ICut^{n}}{\hookrightarrow}\;P$ $\framebox{n}\;\ell\mapsto v$ Strong Alloc and Return to/from stack
Actor (Gordon, 2019) @ι P Variable values Weak Send Message
Memory-Fence (Doko and Vafeiadis, 2016, 2017; Dang et al., 2019) $\triangle_{\pi}$ and $\triangledown_{\pi}$ $\ell\mapsto v$ Weak Fence Acquire and Release
Address Space (Kuru and Gordon, 2025) [r]P $\ell\mapsto v$ Weak Address Space Switch
Ref-Count (Wagner et al., 2024) @ P $\ell_{1}\mapsto v$ Weak Allocating, Dropping and Sharing a Reference
*The StackRegion Modality is an instance of NextGen (called the Independence Modality in (Vindum et al., 2025)).
@@
3.3. Resource
Consider first the address space abstraction in an OS kernel. An address space of a process is a container of virtual addresses referencing data in memory. One would expect to have points-to assertions from separation logic to specify ownership of a memory reference pointing to some data. In the classic setting (Reynolds:2002:SLL:645683.664578), such a points-to assertion $x\mapsto v$ asserts ownership of memory location $x$, and that the contents of that memory location are $v$. But that ownership is relative to a specific address space — a specific container. We tend to think directly about what is true in an address space, with the simplest piece being an association between a virtual address and the data it points to. We call the simplest piece, in this and other examples, the resource element:
Definition 3.0 (Resource Elements).
@@ 3.3. Resource
Definition 3.0 (Resource Context).
@@
Except for the Post-Crash-Modality, one can think of the resource contexts in the first column in Table 1 as containers for the corresponding resource elements in the second column.
@@
When reasoning about stack frame contents, the resource element would be a stack-memory points-to assertion ( $\framebox{n}\;\ell\mapsto v$) indicating that a certain offset into stack region $n$ holds value $v$. The regions (frames) themselves are the contexts.
@@
For virtual memory management, a virtual-points-to, $\ell\mapsto v$, is an ownership assertion pairing a virtual address ($\ell$) with data ($v$) in an address space, which is a resource context for virtual address mappings, $\ell\mapsto v$, and uniquely identified with a root address. Logically, specifying an address space switch  (Kuru and Gordon, 2025) from the address space with the root address $r2$ to the one with the root address $r1$
$\{[r1]P\ast Q\}cr3:=r1\{P\ast[r2]Q\}_{\textsf{cr3}}$
requires having the resource elements loaded into the memory
$P\triangleq\ell_{1}\mapsto v$...
to be in the resource context of the address space modality with the root $r1$, and, once the new address space is loaded—i.e., cr3 is loaded with the other address space root address—the resource elements ($Q$) that were loaded in the previous view of the memory are now introduced to the resource context with the root $r2$.
@@
When considering weak memory models, we consider address-value mappings of thread-local views in C11 memory model (Lahav et al., 2017)). Modalities $\Delta_{\pi}(P)$ and $\nabla_{\pi}(P)$ with $P$ represent the memory address-value mappings that held before or will hold after certain memory fence operations by thread $\pi$. The elements are points-to assertions, each specific to views of memory after those fence operations.
@@
When dealing with reference counting APIs, we may care to specify reachability of memory nodes ($\ell\mapsto v$) in a certain context defined by a shared root address. A shared memory address $\ell$ can be the root of the graph ($_{\ell}\;(\ell_{1}\mapsto v_{1}\;\ast\;...\;\ast\;\ell_{j}\mapsto v_{j})$) that can be a container of memory nodes ($\ell_{1}\mapsto v$) that are reachable from the root $\ell$.
Actors
When specifying the consistency of an actor in the current view, we need to refer to other actors’ state: $_{\iota}(P)$ asserts that $P$ is true in the actor referenced actor ID $\iota$. Assertions may mention local variables and state at the designated actor by name, so the interpretation of variables in assertions depends on which actor the assertion applies to.
@@
The resource element of Post-Crash Modality is not obvious in Table 1, and needs a bit of explanation. Perennial, based on the Iris logic, has both disk-points-to assertions $d[\textsf{p}]\mapsto_{n}v$ (for a specific disk $d$) and in-memory points-to assertions $\ell\mapsto_{n}^{\overline{\gamma}}$. Perennial crash-recovery logic book-keeps resource names (can be thought of as logical variables) $\overline{\gamma}$ to identify which assertions (resource elements) remain valid after a crash — these assertions are only usable while the names in $\overline{\gamma}$ are valid, and a crash resets them, discarding assumptions about volatile state. A subtlety of the notion of a resource context is that, unlike the earlier examples, the context does not need to be a literal data structure. It can instead be (various forms of) a set of executions, as in the Post-Crash and NextGen modalities. The Post-Crash modality $\diamond$ expresses that the assertion $P$ will be true after a crash discards all unstable storage (i.e., RAM). This was the inspiration for the NextGen modality, which is in fact a framework for defining “after-tt” modalities, where $t$ is an transformation of the global state. [^3] The transformations are subject to some technical constraints that are unrelated to our point here.
3.4. Nominals for Systems Resource Contexts
@@
Strong nominalization is the case where resource elements must explicitly include the identity of their intended context, while weak nominalization occurs when the resource elements implicitly pick up the relevant context from how they are used. The first three modalities in Table 1 are strongly nominalized, with the resource elements generally carrying identifiers of a specific modality usage. [^4] The Post-Crash modality does not look like this in the presentation here; technically the definition of $\diamond$ quantifies over names $\overline{\gamma}$ internally, dealing with sets of possible contexts. The rest are weakly nominalized.
@@
By contrast, weakly nominalized elements are more concise, but then make talking about multiple contexts together marginally more complex: changing which modality an assertion is used with drastically changes its meaning. In the case of the Ref-Count modality, $_{\ell}(\ell_{1}\mapsto v)$ says that $\ell$ points to a reference count wrapping $\ell_{1}$, while placing the $\ell_{1}\mapsto v$ under a jump modality for a different location entails talking about a different region of memory.
@@
4. Conclusion
We have argued that there is a systematic pattern emerging in the use of modalities in the verification of low-level systems software, and explained how a range of modern work fits into this pattern. While most of our examples come from public verification work carried out in Iris, we have one example of these ideas applying in Dafny. We are applying these ideas in ongoing work using different modalities for different parts of a copy-on-write filesystem (e.g., for assertions true in different snapshots), and are observing more structured specifications as a result. So not only are the modalities observably present, but they are demonstrably useful for conducting proofs of systems software. We believe that these essential patterns in the designs of modal abstractions for systems verification constitute a fundamental concept when working out how to specify different kinds of systems code.